OpenStudy (anonymous):

log5(base)^x + log5(base)^6x+5 =1 Please help

OpenStudy (ash2326):

Is this your question? \[\log_{5} x+ \log_{5} 6x+5=1\]

OpenStudy (anonymous):

\[\sqrt{x} \sqrt{6x+5}\] should be added

OpenStudy (ash2326):

Is this your question \[\log_{5} {\sqrt x}+\log_{5} {\sqrt{6x+5}}=1\] ?

OpenStudy (anonymous):

yes.

OpenStudy (ash2326):

We know that \[ \log_c a+ \log_c b= \log_c {ab}\] here we have \[ \log_5 {\sqrt x}+\log_5 \sqrt{6x+5}= \log_{5} \sqrt {x\times (6x+5)}\] so we get \[\log_{5} \sqrt {x\times (6x+5)}=1\] also \[\log_{c} a^b=b* \log_{c} {a}\] so \[\frac{1}{2}\log_{5} {x\times (6x+5)}=1\] we get now \[\log_{5} {x\times (6x+5)}=2\] \[x*(6x+5)=25 \] \[6x^2+5x-25=0\] Can you solve now?

OpenStudy (anonymous):

Wow. Yes I can solve from here. I guess I just need to memorized the formulas. We went over this in class last night for the first time. I appreciate your help, thanks.

OpenStudy (ash2326):

Welcome:) Tia1977 you'll get two values for x, your answer is the one which is positive

OpenStudy (anonymous):

ok